Utilizing $10 dollars of inputs Spain could produce $20 worth of wine or $30 worth of Beer.

Utilizing $10 dollars of inputs France could produce $50 worth of wine or $50 worth of Beer.

Spain has comparative advantage in both beer and wine

France has comparative advantage in both beer and wind

France has absolute advantage and comparative advantage in beer

France has absolute advantage and comparative advantage in wine
Answer #1

Option 4

Spain opportunity costs

1 Beer = 20/30=0.67 wine

1 wine = 30/20=1.5 beer

France opportunity costs

1 Beer = 50/50=1 Wine

1 Wine =50/50=1 Beer

So, France has comparative and absolute advantage in wine as it has forgo lesser beer.